# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v3

@@ 6 Discussion
Explainability is essential in medical imaging and neuroimaging fields. The study Mamalakis et al. (2024) categorized the need for AI explanations into self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery, based on the variability of expert opinions, the stability of the evaluation protocol and the representation dimensionality of the application. To that end we apply their proposed guidance in the task of the binary classification task related to secondary brain sulcal region highly related with psycotic patient and for the first time, we develop a new-patterns discovery XAI 3D-Framework.
The XAI 3D-Framework was developed to provide robust and faithful global and local explanations for a deep learning network in a binary classification task related to brain sulcal topology. We focused on the classification task of determining the presence or absence of the paracingulate sulcus and employed two different 3D deep learning networks: a simple deep 3D convolutional neural network and a two-headed attention layer network. The global explanations generated by our framework surpassed those produced by GradCam and SHAP in terms of faithfulness, providing a robust interpretation of the deep networks for the binary classification task.
Our carefully designed and developed XAI 3D-Framework not only offers local and global interpretations but also identifies significant sub-regions involved in the existence and absence of the PCS, a highly variable secondary sulcus region highly correlated with psycotic patient. We discovered significant associations between specific anatomical regions of the brain through grey-white surface and sulcal skeleton inputs, that were distinct in the presence or absence of PCS. These findings enhance our understanding of the neural mechanisms related to the developmental origins of the PCS and provide valuable insights for future research in the field of neurological and psychiatric disorders. The regions of focus, both stemming from the grey/white matter interface and sulcal skeleton inputs, call for a globalized reflection about the neurodevelopmental mechanisms leading to the formation of a paracingulate sulcus (or absence thereof).
The focus which is decentralized from the region of definition of the paracingulate sulcus, suggests that the most relevant information about its presence of the paracingulate sulcus is located elsewhere, which is counter intuitive. Yet, this result may relay developmental insights regarding the mechanisms leading to the potential formation of the paracingulate sulcus De Vareilles et al. (2023). When looking at the sulcal skeleton results, while the timing of apparition of the internal parietal and sub-parietal sulci are unreported, the superior temporal and prefrontal sulci are both reported to first appear around 26 weeks of gestational age in the fetus Garel et al. (2001); Habas et al. (2012), while the paracingulate sulcus (which would be part of the secondary cingulate folding) only starts to appear around 31 weeks of gestational age Garel et al. (2001). The relationship between those sulci could therefore indicate an earlier chronology of the events orchestrating the apparition of a paracingulate sulcus. This could hint towards a greater genetic constraint on this sulcus as earlier developing sulci have been reported to show more heritability Pizzagalli et al. (2020).
Incidentally, the relevance of the Sylvian fissure in the results may imply an intertwining of gyrification and opercularization in the mechanisms leading to the formation of a paracingulate sulcus. This could be a potential explanation for the commonly reported leftward predominance of the left paracingulate sulcus, as, while gyrification shows a rightward developmental advance, opercularization has been reported to show a leftward advance De Vareilles et al. (2023). Therefore, the concurrent intervention of these two phenomenons could justify the asymmetrical patterning of the developing paracingulate sulcus and potentially its resulting form; in that case, the presence of this sulcus could be influenced by even earlier developmental events as opercularization starts before gyration and is importantly genetically regulated Mallela et al. (2020)
@@
Our study introduces an automated and impartial annotation process, enhancing the possibilities for comprehensive explorations of sulcal development in relationship to cognitive and behavioral development. In the realm of deep learning and neuroscience, our contribution is a pioneering 3D explainable framework validating pattern learning in computer vision classifications tasks. Applied to paracingulate detection, our frameworks allowed to explore anatomical covariability and incidentally delve into the chronology of events intertwining structural and functional consequences. The effectiveness of our framework remains dependent on the quality and diversity of accessible data, and its performance could differ in alternate contexts. Additionally, while the incorporation of explainability techniques contributes to interpretability, it might not encompass all underlying intricacies of network behavior. The selection of the two local methods necessitates further exploration of alternative approaches to ensure comprehensiveness. The enhancement of classification outcomes demands the incorporation of additional information beyond structural MRI data to achieve improved accuracy.
@@
To further investigate and incorporate more intricate strategies, the combination of XAI techniques and reduce dimensionality methods can explore overlapping aspects. Furthermore, complex approaches like t-SNE for non-linear dimensionality reduction can be employed to delve deeper into the subject. Moving forward, we envision expanding the framework’s applicability to other neurological conditions and classification tasks (such as shizophrenia or bipolar disorder), employing external datasets like BeneMIN datasets Lisiecka et al. (2015), Deakin et al. (2018), and further refining the interpretability techniques. Our study paves the way for a more systematic exploration of sulcal variability through the lens of deep learning, which can help explore cognitive and functional variability as well as pathological changes.
@@ 7 Conclusion
In this study, we developed a novel XAI 3D-Framework to deliver both local and global explanations for new-patterns discovery task in the secondary sulcus region PCS. The global explanations derived from this framework outperformed established methods such as SHAP and GradCam in terms of faithfulness, achieving the quality needed to discover new global patterns. Additionally, the three-dimensional representation of the proposed framework allows for the robust identification of complex new patterns within the intricate dimensionality of neuroscience applications. To accomplish this, the 3D-Framework enables the identification of sub-regions that are critical to decision-making through a post-fusion transformation of global explanations and statistical features.
@@
This work advances the fields of both deep learning and neuroscience, enabling automated and impartial annotations while illuminating intricate neuroanatomical relationships across diverse subregions. Our XAI 3D-Framework offers unparalleled insights, thus setting the stage for comprehensive exploration within the ever-evolving domains of deep learning and neuroscience.

### v4

@@ 6 Discussion
Explainability is essential in medical imaging. Healthcare applications of AI need to able to explain their decision making to build trust and ensure that their predictions align with other symptoms and signs that affect health. Neuroimaging, the combination of brain images and computational methods, is a research application of medical imaging. Here, explainability for AI predictions support the assessment of the validity of results, but can also identify key contributors to decisions that in themselves reveal new patterns and directions for future investigation.
@@
Our prior study (Mamalakis et al. (2024)) categorized the need for AI explanations into self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery, based on the variability of expert opinions, the stability of the evaluation protocol and the representation dimensionality of the application. We applied the proposed guidance in a binary classification task related to a sub-region of the medial surface of the brain where secondary sulci are highly reproducible related with symptoms of psychosis, specifically hallucinations. This application was of the new-pattern discovery class. The output of the explainability indicated a wide distribution of brain regions on which the predictions depend suggesting covariant development of these regions during the perinatal period (De Vareilles et al. (2023)).
Automatic classification of psychotic and control patients based on structural MRI is a challenging task (Nunes et al. (2020)). In most cases, an acceptable detection rate is around 80.0% in clinical applications. However, in highly heterogeneous and variable cohorts, a lesser performance can be acceptable; approximately 60.0-70.0% acceptance accuracy; Iniesta et al. (2018); Hosmer and Lemeshow (2000). The TOP-OSLO cohort was particularly difficult for classification tasks of psychotic and control patients using structural MRI, with an accuracy of less than 60.0% (Nunes et al. (2020)). The variability of the paracingulate region in the TOP-OSLO cohort and the heterogeneity of the dataset create a highly challenging context for the automatic binary classification task of PCS presence. In this study, an accuracy of more than 70.0% in the left hemisphere and more than 60.0% in the right hemisphere was achieved, delivering an acceptable automated 3D deep learning network (Iniesta et al. (2018); Hosmer and Lemeshow (2000)) to apply global explainability methods for new-pattern discovery (Mamalakis et al. (2024)).
For the binary classification task of PCS presence, we developed two different 3D deep learning networks: a simple 3D convolutional neural network and a two-headed attention layer network. These networks utilized 3D brain inputs derived from preprocessed structural MRI scans, which included grey-white surface boundaries and sulcal skeletons from both hemispheres of a well-annotated cohort of 596 subjects. The performance of all networks was higher in the left hemisphere than in the right hemisphere. This discrepancy in performance was expected as the PCS is more prominent in the left hemisphere, including in psychopathological situations such as schizophrenia (Garrison et al. (2015)). Moreover, the left PCS has more associations with regional cortical thickness and sulcal depth than the right PCS, implying greater covariability of anatomical features in our input modalities with the presence of the PCS in the left hemisphere (Fornito et al. (2006)).
We developed an XAI 3D-Framework to fill the gap in the literature by providing accurate, low-complexity, three-dimensional (3D) global explanations in the neuroimaging field, where complex representational spaces demand more than basic two-dimensional interpretations. This framework was designed to offer robust and faithful global explanations for a deep learning network in the binary classification task of PCS presence. Our global explanations surpassed those produced by GradCam and SHAP in terms of faithfulness, providing a reliable interpretation of the deep networks for this classification task.
The overall explainability outputs cover wide regions of the brain, but we can notice some repetitive patterns through the different pipelines and modalities. In particular, there is a repeat of the cingulate region, the posterior temporal region, and both the medial and inferior frontal cortices. This may reflect some neurodevelopmental intertwining of the macroscopical development of the PCS and these regions. Regarding the cingulate and medial frontal regions, this intertwining is rather self-explanatory as the PCS is located in the medial frontal region, directly adjacent to the cingulate region, and as such the developmental events leading to the formation of a PCS are very likely to affect these regions. The two other regions mostly standing out are the inferior frontal region and the posterior temporal region. In the fetus, the sulci matching these regions (namely the inferior frontal sulcus and the posterior superior temporal sulcus) have both been reported to start appearing at 26 weeks of gestational age (w GA), while the cingulate sulcus appears earlier (around 23w GA) and the “secondary cingular sulci”, which encompass the PCS, appear later, at 31w GA (Garel et al. (2001)). This may point towards a time-window which is decisive to the development of the PCS, anterior to its actual apparition.
In terms of functional interpretation, it is interesting to notice a striking similarity between the regions on which the AI mostly focuses and the regions which have been reported to show the most functional connectivity with the paracingulate region (Fedeli et al. (2022); Harper et al. ), and the anterior cingulate region (Lopez-Persem et al. ). These studies report relevant functional connectivity between the medial frontal lobe (including the PCS) and the temporal region (including a focus on the posterior superior temporal region), the inferior frontal region, and the medial parietal region, which are all regions showing particular interest in the present work. Both Fedeli et al. (2022) and Harper et al. investigate the relation between the presence of a PCS and the related functional connectivity in these regions, and obtain more focused results (respectively in the cerebellum and superior anterior temporal region, or in the medial frontal region), but the important functional relationships between these regions support the relevance of the regions highlighted by our results.
@@
The effectiveness of our framework remains dependent on the quality and diversity of accessible data, and its performance could differ in alternate contexts. Additionally, while the incorporation of explainability techniques contributes to interpretability, it might not encompass all underlying intricacies of network behavior. The selection of the two local methods necessitates further exploration of alternative approaches to ensure comprehensiveness. The enhancement of classification outcomes demands the incorporation of additional information beyond structural MRI data to achieve improved accuracy. To further investigate and incorporate more intricate strategies, the combination of XAI techniques and reduce dimensionality methods can explore overlapping aspects. Furthermore, complex approaches like t-SNE for non-linear dimensionality reduction can be employed to delve deeper into the subject.
@@
Moving forward, we envision expanding the framework’s applicability to other neurological conditions and classification tasks (such as schizophrenia or bipolar disorder), employing other external datasets and further refining the interpretability techniques. Our study paves the way for a more systematic exploration of sulcal variability through the lens of deep learning, which can help explore cognitive and functional variability as well as pathological changes.
@@ 7 Conclusion
In this study, we classified the presence and absence of the secondary sulcus, the PCS, in a highly heterogeneous, unbiased, and well-annotated cohort of 596 structural MRIs. We used traditional 3D deep convolutional neural and attention mechanism networks to achieve acceptable detection accuracy and apply explainability techniques for new-pattern discovery.
As there were no global explainability methodologies in the literature efficient enough to deliver robust, accurate, low-complexity and three-dimensional explanations, we developed and proposed a novel XAI 3D-Framework to provide global explanations for the task of PCS detection. The global explanations derived from this framework outperformed established methods such as SHAP and GradCam in terms of faithfulness, achieving the quality needed to discover new global patterns. Additionally, the 3D representation of the proposed framework allows for the robust identification of complex new patterns within the high complexity and dimensionality of neuroscience applications. To accomplish this, the 3D-Framework enables the identification of sub-regions that are critical to decision-making through a post-fusion transformation of global explanations and statistical features.
The detection of the presence of the PCS appeared to focus most reliably on features related to the posterior temporal region and the internal parietal region (on the sulcal skeleton), with additional focus on the cingulate region and thalamus when focusing on the grey-white surface. This highlights potential co-variation between these structures which should be investigated further, notably in terms of common genetic influences in these regions, both for neurodevelopmental and functional or pathological considerations.
@@
This work advances the fields of both deep learning and neuroscience, enabling automated and impartial annotations while illuminating detailed neuroanatomical relationships across diverse subregions. Our XAI 3D-Framework offers unparalleled insights, thus setting the stage for comprehensive exploration within the ever-evolving domains of deep learning and neuroscience.